OpenStudy (anonymous):

Five loaves of bread are to be cut into pieces that are each 1/3 the size of a whole loaf. How many pieces will there be?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

do you know how i got that?

OpenStudy (anonymous):

You have five loaves of bread and you're making each piece 1/3 the side of one loaf, so: \[\frac{ 5 }{ \frac{ 1 }{ 3 } } = 5 \times 3 = 15 pieces\]